# Recurrent Complex-Weighted Autoencoders for Unsupervised Object Discovery

### v2

@@ Abstract
Current state-of-the-art synchrony-based models encode object bindings with complex-valued activations and compute with real-valued weights in feedforward architectures. We argue for the computational advantages of a recurrent architecture with complex-valued weights. We propose a fully convolutional autoencoder, SynCx, that performs iterative constraint satisfaction: at each iteration, a hidden layer bottleneck encodes statistically regular configurations of features in particular phase relationships; over iterations, local constraints propagate and the model converges to a globally consistent configuration of phase assignments. Binding is achieved simply by the matrix-vector product operation between complex-valued weights and activations, without the need for additional mechanisms that have been incorporated into current synchrony-based models. SynCx outperforms or is strongly competitive with current models for unsupervised object discovery. SynCx also avoids certain systematic grouping errors of current models, such as the inability to separate similarly colored objects without additional supervision.
@@ 1 Introduction
Consider the shapes composed of horizontal and vertical bars in Figure 1: the letters $\mathbb{T}$ and $\mathbb{H}$ and a pair of overlapping squares with occlusion. Perceptual grouping involves determining which of the bars belong together. The pairs highlighted in green are part of the same object; the pairs highlighted in red are parts of different objects. A feedforward convolutional architecture is not well suited to this task because spatially local patches are ambiguous with regard to grouping. While a feedforward, fully connected architectures may work in principle, it is also problematic because statistical regularities in images are local and fairly low order (i.e., involve a subset of image features). However, if we drop the feedforward restriction, a convolutional model can iteratively converge on a solution: each iteration suggests soft constraints on how features should be grouped, and over iterations, constraints can propagate from one region in the image to adjacent regions. This methodology was adopted in classical AI vision models [28] and in early synchrony-based neural networks [20]. In the latter, phases—representing object binding—are initialized to random values and each iteration, the phases are updated to be consistent with constraints from neighboring patches. After multiple iterations the model converges on an interpretation that assigns each image location to a particular phase (object).
@@
Critical to this account is the use of complex-valued weights. Just as hidden units in a standard architecture detect configurations of features, hidden units in a complex-valued analogue detect configurations of features in a particular relative phase arrangement. A hidden unit that is tuned to the top of a $\mathbb{T}$ will expect the horizontal and vertical bar to have the same phase; a hidden unit that is tuned to occlusion (the red patches in Figure 1) will expect the horizontal and vertical bars to be out of phase. In this manner, the hidden units can encode alternative hypotheses concerning object groupings, and iterative processing will attempt to find global configurations that are locally consistent.
Figure 2: SynCx is a fully convolutional autoencoder that iteratively processes the input image. It starts with a randomly initialized phase $\bm{\phi}_{x}^{1}$ and updates the phases in a stateful manner, i.e., output phase at iteration 1 fed as input at iteration 2 ($\bm{\phi}_{x}^{2}\leftarrow\bm{\phi}_{z}^{1}$) and so on.
@@
Drawing on these intuitions we design a model and training strategy for the fully unsupervised setting. Figure 2 shows our proposed model, Synchronous Complex Network (SynCx), which uses complex-valued weights (detectors) to process complex-valued activations (features) in every layer of a fully convolutional autoencoder. The matrix vector product between complex-valued weights and activations ensures that weights process inputs not only based on their features (magnitudes $\rightarrow$ features) but also their phase relationships (phases $\rightarrow$ object bindings). The addition and multiplication operators in complex algebra mechanistically implement the binding mechanism in our model removing the need for additional gating mechanisms ($\chi$ [22] or cosine binding [27]) or contrastive training [24]. The spatial locality of constraints are ensured via 2D convolutions used at every layer in the autoencoder. The model is initialized with a random phase map (prior on object bindings) and learns to (de)synchronize groups of features based on their (dis)similarity iteratively. The autencoder weights are shared across iterations and predicted output phases (top-down feedback) from the previous iteration are used as the input phases at the next iteration while the pixel values are the input magnitudes at each step. Iteratively updating the phases allows spatially local constraints to be slowly propagated across the spatial axes and in principle should lead to the system settling to a fixed point. The autoencoder is trained to simply reconstruct the input image at each iteration.
@@ 2 Method
Notice in Equation 3, that the magnitude components are always clamped to the image ($\mathbf{x}^{\prime}$) while the phase components ($\bm{\phi}_{x}^{n}$) are fed back. At the $n^{\textrm{th}}$ iteration, $\mathbf{x}^{n}$ is processed by the encoder and decoder layers, denoted as Net$(\mathbf{x}^{n})$, to compute a complex-valued output $\mathbf{z}^{n}\in\mathbb{C}^{h\times w\times c}$. The magnitude component of $\mathbf{z}^{n}$, i.e. $\bm{\mu}_{z}^{n}$, is the reconstruction of input image $\bm{\mu}_{x}$. The phase component of $\mathbf{z}^{n}$, i.e. $\bm{\phi}_{z}^{n}$, is used to initialize the complex-valued input $\mathbf{x}^{n+1}$ at the next iteration:
@@
Here, we outline the processing steps used to extract discrete object assignments for every pixel from its continuous phase values. We start by taking the output complex-valued feature map $\mathbf{h}^{N}\in\mathbb{C}^{h\times w\times d_{\text{out}}}$ from an appropriate layer (typically penultimate) of the decoder at the last iteration $N$. We use latent-level complex-valued features instead of output-level so as to use high-order features (texture/shape/color etc.) than simply color cues (RGB space) to determine their object constituency. Then, we construct a complex-valued feature map with unit magnitudes (i.e., $\bm{\mu}_{h}^{N}\leftarrow\mathbf{1}\in\mathbb{R}^{h\times w\times d_{\text{out}}}$) and phases the same as that of $\mathbf{h}^{N}$ (i.e., $\bm{\phi}_{h}^{N}$). Background regions in this constructed feature map are masked out by setting their magnitudes to zero and then converted to the Cartesian form element-wise. Finally, these features in Cartesian form are clustered using $k$-Means to compute the object assignments for every location in the spatial map. Using unit magnitudes for complex-valued activations ensures that object assignments are solely determined by their orientations. For more details on the object assignment process refer to Appendix A.
@@ 3 Results
We seek to qualitatively inspect phase components of the output complex-valued feature map $\mathbf{h}\in\mathbb{C}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$ from an intermediate decoder layer to see if our model has indeed learned phases specialized in an object-centric manner. We visualize the phase component of the high-dimensional feature map as a heatmap in two dimensions. We assign unit magnitudes and retain the phase components of $\mathbf{h}$ and elementwise convert these to the cartesian form as we did previously to extract object assignments. We apply t-SNE [38] to perform dimensionality reduction on the complex-valued feature map of an input image to recover a scalar ‘composite’ phase value at every spatial location (see Appendix A for more details). This ‘composite’ phase map is visualized as a two dimensional heatmap and radial plot with colors in the former corresponding to orientations in the latter (see columns 2 & 3 in Figure 3). From Figure 3 shows how the phase maps evolve across the iterations in SynCx grouping process. In the sample image from Tetrominoes (row 1 in Figure 3), we can see how the phase ‘bands’ corresponding to each tetris block are progressively separated in their orientations. Such a visualization gives an interpretable artefact that allows a qualitative inspection of the phase synchronization process.
@@
Table 1 compares the performance of recent state-of-the-art synchrony-based baselines (CAE, CAE++, CtCAE, and RF) against ours (SynCx) on the unsupervised object discovery task. On Tetrominoes, we observe that SynCx outperforms all baselines on grouping performance. SynCx more gracefully separates objects of the same color (see Figure 4) compared to RF which in addition to complex-valued activations requires $\chi$-binding mechanism. SynCx also outperforms the CtCAE baseline which in addition to complex-valued activations requires $\chi$-binding and contrastive training to separate similarly colored objects. On dSprites, SynCx strongly outperforms CAE, CAE++ and CtCAE baselines while being strongly competitive with RF. Similarly on CLEVR, SynCx strongly outperforms CAE, CAE++ and CtCAE baselines while being competitive with RF. Overall, SynCx despite its simple binding mechanism (no gates) and training strategy (no contrastive training) outperforms or strongly competitive with all recent state-of-the-art synchrony-based models. We refer to Appendix B for the model comparisons that includes the MSE loss as well. We refer to Appendix C for additional qualitative grouping examples from our model.
Table 1: ARI scores (mean $\pm$ standard deviation across 5 seeds) for CAE, CAE++, CtCAE, RF and SynCx models for Tetrominoes, dSprites and CLEVR. Results for CAE, CAE++ and CtCAE baselines are taken from Stanić et al. [24].
@@
We measure the effect of having representational bottlenecks in the hidden layers of our model. To do so, we design a variant of our model (denoted as SynCx w/o bottleneck in Table 2) that preserves the spatial resolution of feature maps in the encoder (using stride of 1 in all layers) thereby removing all bottlenecks.
@@
We observe a sharp drop in grouping performance for our ablated model variant without any bottlenecks despite it achieving a lower test loss (see Figure 5). This result supports our initial intuition that representational bottlenecks force an autoencoder with complex-valued activations to compress spatial regularities in features by using phase components to capture relationships between them. Since objects are modular units containing highly regular features within their boundaries, this leads to phases strongly specializing (synchronizing) towards them to facilitate better image compression.
@@
Table 3: No. of iterations ablation on dSprites.
@@
We measure the effect of multiple iterations to update phases on the grouping performance of our model. Table 3 shows the grouping performance of our model with increasing number of iterations. We observe a general trend of increasing phase specialization towards objects and better reconstructions as we increase the number of iterations. This supports our qualitative characterization of the models’ operation — starting with random object assignments at each iteration it updates the hypotheses about object assignment for every pixel.
@@
We measure the effect that the random initialization of phases has on the grouping performance of our model. To achieve this, we compare (see Table 4) the grouping performance of one variant which samples the initial phases from a uniform distribution between -$\pi$ and $\pi$ versus the other which samples from a von-Mises distribution with mean of 00 and concentration of $1$.
@@
Rest of the model and training hyperparameters are kept the same between these two variants. We observe that the variant using von-Mises distribution to sample initial input phases achieves better test loss and improved grouping scores. Therefore, we use it as the default initialization for input phases in all experiments.
@@
We test how reliant on color cues are the two best performing synchrony-based models from Table 1, i.e., RF and SynCx, for unsupervised object discovery. We train and evaluate them on a grayscale version of the CLEVR dataset.
@@
To separate objects in this dataset, the models’ would need to rely more on other features like shape, texture, etc. From Table 5, we see that the grouping performance of RF on grayscale CLEVR drops sharply compared to the original (see Table 1) indicating its heavy reliance on color cues. Our models’ grouping performance also drops but to a much lesser degree compared to RF. This indicates a weaker dependence on color cues in the binding mechanism of our model. In fact, on grayscale CLEVR SynCx significantly outperforms RF despite having slightly worse performance on the original colored version. The binding mechanism (i.e., matrix-vector product) in our model updates phases based on agreement between feature detectors (complex-valued weights) and high-dimensional attributes (complex-valued activations) that model color, edges, texture, shape etc. while taking it account their phase relationships (local context). This facilitates it to bind to objects more robustly and not simply rely on color. Overall, we can see that despite its simple architecture and training procedure our model shows good grouping performance on a wider range of visual environments. However, binding mechanisms in state-of-the-art synchrony models are still not quite robust. They are unable to reliably capture objects even on synthetic datasets like CLEVR containing simple 3D shapes with metallic or matte textures under camera lighting with a non-textured background.

### v3

@@ Abstract
Current state-of-the-art synchrony-based models encode object bindings with complex-valued activations and compute with real-valued weights in feedforward architectures. We argue for the computational advantages of a recurrent architecture with complex-valued weights. We propose a fully convolutional autoencoder, SynCx, that performs iterative constraint satisfaction: at each iteration, a hidden layer bottleneck encodes statistically regular configurations of features in particular phase relationships; over iterations, local constraints propagate and the model converges to a globally consistent configuration of phase assignments. Binding is achieved simply by the matrix-vector product operation between complex-valued weights and activations, without the need for additional mechanisms that have been incorporated into current synchrony-based models. SynCx outperforms or is strongly competitive with current models for unsupervised object discovery. SynCx also avoids certain systematic grouping errors of current models, such as the inability to separate similarly colored objects without additional supervision.  [^1] Official code repository: https://github.com/agopal42/syncx
@@ 1 Introduction
Consider the shapes composed of horizontal and vertical bars in Figure 1: the letters $\mathbb{T}$ and $\mathbb{H}$ and a pair of overlapping squares with occlusion. Perceptual grouping involves determining which of the bars belong together. The pairs highlighted in green are part of the same object; the pairs highlighted in red are parts of different objects. A feedforward convolutional architecture is not well suited to this task because spatially local patches are ambiguous with regard to grouping. While a feedforward, fully connected architectures may work in principle, it is also problematic because statistical regularities in images are local and fairly low order (i.e., involve a subset of image features). However, if we drop the feedforward restriction, a convolutional model can iteratively converge on a solution: each iteration suggests soft constraints on how features should be grouped, and over iterations, constraints can propagate from one region in the image to adjacent regions. This methodology was adopted in classical AI vision models [28] and in early synchrony-based models [20]. In the latter, phases—representing object binding—are initialized to random values and each iteration, the phases are updated to be consistent with constraints from neighboring patches. After multiple iterations the model converges on an interpretation that assigns each image location to a particular phase (object).
@@
Complex weights are a natural choice to instantiate this function in our synchrony formulation outlined above. Just as hidden units in a standard architecture detect configurations of features, hidden units in a complex-valued analogue detect configurations of features in a particular relative phase arrangement. A hidden unit that is tuned to the top of a $\mathbb{T}$ will expect the horizontal and vertical bar to have the same phase; a hidden unit that is tuned to occlusion (the red patches in Figure 1) will expect the horizontal and vertical bars to be out of phase. In this manner, the hidden units can encode alternative hypotheses concerning object groupings, and iterative processing will attempt to find global configurations that are locally consistent.
@@
Drawing on these intuitions we design a model and training strategy for the fully unsupervised setting. Figure 2 shows our proposed model, Synchronous Complex Network (SynCx), which uses complex-valued weights (detectors) to process complex-valued activations (features) in every layer of a fully convolutional autoencoder. The matrix vector product between complex-valued weights and activations ensures that weights process inputs not only based on their features (magnitudes $\rightarrow$ features) but also their phase relationships (phases $\rightarrow$ object bindings). The addition and multiplication operators in complex algebra mechanistically implement the binding mechanism in our model removing the need for additional gating mechanisms ($\chi$ [22] / cosine binding [27]) or contrastive training [24]. The spatial locality of constraints are ensured via 2D convolutions used at every layer in the autoencoder. The model is initialized with a random phase map (prior on object bindings) and learns to (de)synchronize groups of features based on their (dis)similarity iteratively. The autoencoder weights are shared across iterations and predicted output phases (top-down feedback) from the previous iteration are used as the input phases at the next iteration while the pixel values are the input magnitudes at each step. Iteratively updating the phases allows spatially local constraints to be slowly propagated across the spatial axes and in principle should lead to the system settling to a fixed point. The autoencoder is trained to simply reconstruct the input image at each iteration.
Figure 2: SynCx is a fully convolutional autoencoder that iteratively processes an input image. It starts with a randomly initialized phase $\bm{\phi}_{x}^{1}$ and the input image $\bm{\mu}_{x}$ in the magnitude-component updates the phases in a stateful manner, i.e., output phase at iteration 1 fed as input at iteration 2 ($\bm{\phi}_{x}^{2}\leftarrow\bm{\phi}_{z}^{1}$) and so on. The magnitude-component at the input is always clamped to the input image $\bm{\mu}_{x}$. SynCx is trained to reconstruct $\bm{\mu}_{x}$ using the output magnitude-component $\bm{\mu}_{z}^{n}$ at every step.
@@ 2 Method
Notice in Equation 3, that the magnitude components are always clamped to the image ($\bm{\mu}_{x}$) while the phase components ($\bm{\phi}_{x}^{n}$) are fed back. At the $n^{\textrm{th}}$ iteration, $\mathbf{x}^{n}$ is processed by the encoder and decoder layers, denoted as Net$(\mathbf{x}^{n})$, to compute a complex-valued output $\mathbf{z}^{n}\in\mathbb{C}^{h\times w\times c}$. The magnitude component of $\mathbf{z}^{n}$, i.e. $\bm{\mu}_{z}^{n}$, is the reconstruction of input image $\bm{\mu}_{x}$. The phase component of $\mathbf{z}^{n}$, i.e. $\bm{\phi}_{z}^{n}$, is used to initialize the complex-valued input $\mathbf{x}^{n+1}$ at the next iteration:
@@
Here, we outline the processing steps used to extract discrete object assignments for every pixel given its continuous phase values. We start by taking the output complex-valued feature map $\mathbf{h}^{N}\in\mathbb{C}^{h\times w\times d_{\text{out}}}$ from the penultimate decoder layer at the last iteration $N$. We use latent-level complex-valued features instead of output-level so as to use high-order features (texture/shape/color etc.) than simply color cues (RGB space) to determine their object constituency. Then, we construct a complex-valued feature map with unit magnitudes (i.e., $\bm{\mu}_{h}^{N}\leftarrow\mathbf{1}\in\mathbb{R}^{h\times w\times d_{\text{out}}}$) and phases the same as that of $\mathbf{h}^{N}$ (i.e., $\bm{\phi}_{h}^{N}$). Background regions in this constructed feature map are masked out by setting their magnitudes to zero and then converted to the Cartesian form element-wise. Finally, these features in Cartesian form are clustered using $k$-Means to compute the object assignments for every location in the spatial map. Using unit magnitudes for complex-valued activations ensures that object assignments are solely determined by their orientations. We note that it would be most natural to cluster the phases from the output layer at each location to assign it to an object. However, we observed that the object separation shown in the output phases is weaker than that from the high-dimensional latent layer. Therefore, for all results presented here we choose to cluster the higher dimensional latent representations (from the penultimate decoder layer) by location. For more details on the object assignment process refer to Appendix A.
@@ 3 Results
We seek to qualitatively inspect phase components of the output complex-valued feature map $\mathbf{h}\in\mathbb{C}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$ from an intermediate decoder layer to see if our model has indeed learned phases specialized in an object-centric manner. We visualize the phase component of the high-dimensional feature map as a heatmap in two dimensions. We assign unit magnitudes and retain the phase components of $\mathbf{h}$ and elementwise convert these to the cartesian form as we did previously to extract object assignments. We apply t-SNE [39] to perform dimensionality reduction on the complex-valued feature map of an input image to recover a scalar ‘composite’ phase value at every spatial location (see Appendix A and Figure 10 for more details). This ‘composite’ phase map is visualized as a two dimensional heatmap and radial plot with colors in the former corresponding to orientations in the latter (see columns 2 & 3 in Figure 3). From Figure 3 shows how the phase maps evolve across the iterations in SynCx grouping process. In the sample image from Tetrominoes (row 1 in Figure 3), we can see how the phase ‘bands’ corresponding to each tetris block are progressively separated in their orientations. Such a visualization gives an interpretable artefact that allows a qualitative inspection of the phase synchronization process.
@@
Table 1 compares the performance of recent state-of-the-art synchrony-based baselines (CAE, CAE++, CtCAE, and RF) against ours (SynCx) on the unsupervised object discovery task. On Tetrominoes, we observe that SynCx outperforms all baselines on grouping performance. SynCx more gracefully separates objects of the same color (see Figure 4) compared to RF which in addition to complex-valued activations requires $\chi$-binding mechanism. SynCx also outperforms the CtCAE baseline which in addition to complex-valued activations requires $\chi$-binding and contrastive training to separate similarly colored objects. On dSprites, SynCx significantly outperforms CAE, CAE++ and CtCAE baselines while being strongly competitive with RF. Similarly on CLEVR, SynCx strongly outperforms CAE, CAE++ and CtCAE baselines while being competitive with RF. Overall, SynCx despite its simple binding mechanism (no gates) and training strategy (no contrastive training) outperforms or is strongly competitive with all recent state-of-the-art synchrony-based models. However, there still remains a significant gap in grouping performance between synchrony-based models w.r.t dominant slot-based approaches such as SlotAttention [34] especially on CLEVR. We refer to Appendix B for the model comparisons that includes the MSE loss as well. We refer to Appendix C for additional qualitative grouping examples from our model.
Table 1: ARI scores (mean $\pm$ standard deviation across 5 seeds) for CAE, CAE++, CtCAE, RF, SynCx and SlotAttention on Tetrominoes, dSprites and CLEVR. Results for CAE, CAE++ and CtCAE baselines taken from Stanić et al. [24] and for SlotAttention taken from Locatello et al. [34].
@@
We measure the effect of having representational bottlenecks (i.e. spatial resolution of feature maps) in the hidden layers of our model. To do so, we design a variant of our model (denoted as SynCx w/o bottleneck in Table 2) that preserves the spatial resolution of feature maps in the encoder (using stride of 1 in all layers) thereby removing all bottlenecks.
@@
We ensure that both these variants (SynCx and SynCx w/o bottleneck) have the same number of parameters. Instead we only vary the spatial resolution of feature maps, so the bottleneck is w.r.t spatial resolution and not the number of model parameters. We observe a sharp drop in grouping performance for our ablated model variant without any bottlenecks despite it achieving a lower test loss (see Figure 5). This result supports our initial intuition that representational bottlenecks force an autoencoder with complex-valued activations to compress spatial regularities in features by using phase components to capture relationships between them. Since objects are modular units containing highly regular features within their boundaries, this leads to phases strongly specializing (synchronizing) towards them to facilitate better image compression.
@@
Table 3: Ablation on number of iterations used for training on dSprites.
@@
We measure the effect of multiple iterations to update phases on the grouping performance of our model. Table 3 shows the grouping performance of our model with increasing number of iterations used during training. We observe a general trend of increasing phase specialization towards objects and better reconstructions as we increase the number of iterations. This supports our qualitative characterization of the models’ operation — starting with random object assignments it iteratively refines the hypotheses for the assignment for each pixel.
Table 4: Ablation on number of iterations used at test-time on dSprites.
No. of Iterations MSE $\downarrow$ ARI $\uparrow$
4 1.71e-3 $\pm$ 1.38e-4 0.81 $\pm$ 0.01
5 1.71e-3 $\pm$ 1.40e-4 0.82 $\pm$ 0.01
6 1.70e-3 $\pm$ 1.41e-4 0.82 $\pm$ 0.01
We also measure the effect of increasing the number of iterations at test-time compared to that used during training. Table 4 shows the grouping performance of our model as we increase the number of iterations from $4$ to $6$ while during training the model used $3$ iterations. We observe no drop in object separation performance of our model when we extrapolate the number of iterations used at test-time.
@@
We measure the effect that the initialization of phases has on the grouping performance of our model. To achieve this, we compare (see Table 5) the grouping performance of three variants — i) initial phases sampled from a uniform distribution between -$\pi$ and $\pi$ ii) initial phases sampled from a von-Mises distribution with mean of 00 and concentration of $1$ and iii) all values in the initial phase map are set to zero.
@@
Rest of the model and training hyperparameters are kept the same between these two variants. We observe that the variant using von-Mises distribution to sample initial input phases achieves the best test loss and improved grouping scores among all variants. This suggests the variance of the noise distribution is an important factor to tune for phase synchronization. The von-Mises distribution with a mean of 00 and concentration of $1$ is the circular analogue of the Normal distribution and therefore samples “noisy” phase values with an intermediate level of variance compared to the other two alternatives which have maximum and zero variance. We use the von-Mises variant as the default phase initialization for input phases in all experiments.
Table 6: Comparison of parameter counts (rounded up to the nearest thousand) of various synchrony-based models. Total number of parameters expressed in terms of number of real-valued floats.
Model Tetrominoes dSprites CLEVR
CAE++ 5,372,000 11,713,000 22,281,000
CtCAE 5,372,000 11,713,000 22,281,000
RF 6,630,000 11,861,000 22,592,000
SynCx 966,000 974,000 974,000
Table 7: Wall-clock training times for SynCx and RF models on a P100 GPU.
Model Tetrominoes dSprites
RF 7h 20m 4h 50m
SynCx 1h 35m 4h 10m
Parameter and Training Efficiency
Table 6 shows the parameters counts for the various synchrony-based models. We see that our model, SynCx, is 6-23x more parameter efficient compared to the other synchrony-based alternatives. Table 7 shows the wall-clock training times for convergence of SynCx against the strongest synchrony baseline, RF, on Tetrominoes and dSprites. We see that our model on these datasets consistently takes lesser wall-clock time for training in these settings. These results highlight the computational benefits of the simple ideas of complex-valued weights and recurrence that are the essence of our model design.
@@
We test how reliant on color cues are the two best performing synchrony-based models from Table 1, i.e., RF and SynCx, for unsupervised object discovery. We train and evaluate them on a grayscale version of the CLEVR dataset. To separate objects in this dataset, the models’ would need to rely more on other features like shape, texture, etc. From Table 8, we see that the grouping performance of RF on grayscale CLEVR drops sharply compared to the original (see Table 1) indicating its heavy reliance on color cues. Our models’ grouping performance also drops but to a much lesser degree compared to RF. This indicates a weaker dependence on color cues in the binding mechanism of our model. In fact, on grayscale CLEVR SynCx significantly outperforms RF despite having slightly worse performance on the original colored version.
@@
The binding mechanism (i.e., matrix-vector product) in our model updates phases based on agreement between feature detectors (complex-valued weights) and high-dimensional attributes (complex-valued activations) that model color, edges, texture, shape etc. while taking it account their phase relationships (local context). This facilitates it to bind to objects more robustly and not simply rely on color. Overall, we can see that despite its simple architecture and training procedure our model shows good grouping performance on a wider range of visual environments. However, binding mechanisms in state-of-the-art synchrony models are far from perfect. They are unable to reliably capture objects even on synthetic datasets like CLEVR containing simple 3D shapes with metallic or matte textures under camera lighting with a non-textured background.